# DCNv3: Towards Next Generation Deep Cross Network for Click-Through Rate Prediction

### v6

@@ 3.3. Tri-BCE Loss
Tri-BCE loss calculation and trade-off method are shown in Figure 4. We use the widely adopted binary cross-entropy loss (Wang et al., 2023a; Zhu et al., 2023b; Mao et al., 2023) (i.e., Logloss) as both the primary and auxiliary loss for the DCNv3:
@@
$\mathcal{L}_{\scriptscriptstyle D}$ $=-\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}\log\left(\hat{y}_{\scriptscriptstyle D,i}\right)+\left(1-y_{i}\right)\log\left(1-\hat{y}_{\scriptscriptstyle D,i}\right)\right),$
$\mathcal{L}_{\scriptscriptstyle S}$ $=-\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}\log\left(\hat{y}_{\scriptscriptstyle S,i}\right)+\left(1-y_{i}\right)\log\left(1-\hat{y}_{\scriptscriptstyle S,i}\right)\right),$
@@
where $y$ denotes the true labels, $N$ denotes the batch size, $\mathcal{L}_{\scriptscriptstyle D}$ and $\mathcal{L}_{\scriptscriptstyle S}$ represent the auxiliary losses for the prediction results of ECN and LCN, respectively, and $\mathcal{L}$ represents the primary loss. To provide each sub-network with suitable supervision signals, we assign them adaptive weights, $\mathbf{w}_{\scriptscriptstyle D}=\texttt{max}(0,\mathcal{L}_{\scriptscriptstyle D}-\mathcal{L})$ and $\mathbf{w}_{\scriptscriptstyle S}=\texttt{max}(0,\mathcal{L}_{\scriptscriptstyle S}-\mathcal{L})$, and jointly train them to achieve Tri-BCE loss:
(7) $\mathcal{L}_{\text{Tri}}$ $=\mathcal{L}+\mathbf{w}_{\scriptscriptstyle D}\cdot\mathcal{L}_{\scriptscriptstyle D}+\mathbf{w}_{\scriptscriptstyle S}\cdot\mathcal{L}_{\scriptscriptstyle S},$
@@
as demonstrated by (Li et al., 2024b), providing a single supervision signal to sub-networks is often suboptimal. Our proposed Tri-BCE loss helps sub-networks learn better parameters by providing adaptive weights that change throughout the learning process. Theoretically, we can derive the gradients obtained by $\hat{y}_{\scriptscriptstyle D}$:
@@
(8) $\nabla_{(\hat{y}^{+}_{\scriptscriptstyle D})}\mathcal{L}_{\text{Tri}}$ $=-\frac{1}{N}\cdot\frac{\partial\left(\log\hat{y}^{+}+\mathbf{w}_{\scriptscriptstyle D}\log\hat{y}^{+}_{\scriptscriptstyle D}\right)}{\partial\hat{y}^{+}_{\scriptscriptstyle D}}=-\frac{1}{N}\left(\frac{1}{2\hat{y}^{+}}+\frac{\mathbf{w}_{\scriptscriptstyle D}}{\hat{y}^{+}_{\scriptscriptstyle D}}\right),$
@@
$\nabla_{(\hat{y}^{-}_{\scriptscriptstyle D})}\mathcal{L}_{\text{Tri}}$ $=-\frac{1}{N}\cdot\frac{\partial\left(\log(1-\hat{y}^{-})+\mathbf{w}_{\scriptscriptstyle D}\log(1-\hat{y}^{-}_{\scriptscriptstyle D})\right)}{\partial\hat{y}^{-}_{\scriptscriptstyle D}}$
$=\frac{1}{N}\left(\frac{1}{2(1-\hat{y}^{-})}+\frac{\mathbf{w}_{\scriptscriptstyle D}}{1-\hat{y}^{-}_{\scriptscriptstyle D}}\right),$
@@
where $\nabla_{(\hat{y}^{+}_{\scriptscriptstyle D})}$ and $\nabla_{(\hat{y}^{-}_{\scriptscriptstyle D})}$ represent the gradients received by $\hat{y}_{\scriptscriptstyle D}$ for positive and negative samples, respectively. Similarly, the gradient signals received by $\hat{y}_{\scriptscriptstyle S}$ are consistent with those of $\hat{y}_{\scriptscriptstyle D}$, so we do not elaborate further. It can be observed that $\hat{y}_{\scriptscriptstyle D}$ and $\hat{y}_{\scriptscriptstyle S}$ both have the same gradient terms $\frac{1}{2\hat{y}^{+}}$ and $\frac{1}{2(1-\hat{y}^{-})}$, indicating that training both sub-networks with a single loss results in same supervision signals, which is detrimental to the model’s learning. However, our Tri-BCE loss additionally provides dynamically adjusted gradient terms based on $\mathbf{w}_{\scriptscriptstyle D}$ and $\mathbf{w}_{\scriptscriptstyle S}$, ensuring that the sub-networks are directly influenced by the true labels $y$ and adaptively adjust their weights according to the difference between the primary and auxiliary losses. Therefore, Tri-BCE loss provides the sub-networks with more suitable supervision signals.
@@ 3.4. Complexity Analysis
Model Embedding Implicit interaction Explicit interaction Loss Function
DCNv1 (Wang et al., 2017) O($dfs$) O($|W_{\Psi}|$) O($2DL$) O($N$)
DCNv2 (Wang et al., 2021b) O($dfs$) O($|W_{\Psi}|$) O($D^{2}L$) O($N$)
EDCN (Chen et al., 2021) O($dfs$) O($D^{2}L$) O($D^{2}L$) O($N$)
GDCN (Wang et al., 2023a) O($dfs$) O($|W_{\Psi}|$) O($2D^{2}L$) O($N$)
ECN O($dfs$) - O($\frac{D^{2}L}{2}$) O($N$)
[…]
@@
  • Except for our proposed ECN and DCNv3, all other models include implicit interaction to enhance predictive performance, which incurs additional computational costs.
@@
  • Since our DCNv3 uses the Tri-BCE loss, the time complexity of loss computation for DCNv3 is three times that of other models. However, this does not affect the model’s inference speed.
@@ 4. Experiments
In this section, we conduct comprehensive experiments on six CTR prediction datasets to validate the effectiveness, efficiency, and interpretability of DCNv3, and address the following research questions (RQs):
  • RQ1 Do DCNv3 outperform other CTR models in terms of performance? Do they perform well on large-scale and highly sparse datasets?
  • RQ2 Are DCNv3 more efficient compared to other CTR models?
  • RQ3 Do DCNv3 possess interpretability and the ability to filter noise?
@@ 4.1. Experiment Setup
We evaluate DCNv3 on six CTR prediction datasets: Avazu [^1] https://www.kaggle.com/c/avazu-ctr-prediction (Zhu et al., 2021), Criteo [^2] https://www.kaggle.com/c/criteo-display-ad-challenge (Zhu et al., 2021), ML-1M [^3] https://grouplens.org/datasets/movielens (Song et al., 2019), KDD12444https://www.kaggle.com/c/kddcup2012-track2 (Song et al., 2019), iPinYou [^5] https://contest.ipinyou.com/ (Qu et al., 2018), and KKBox [^6] https://www.kkbox.com/intl (Zhu et al., 2022a). Table 2 provides detailed information about these datasets. A more detailed description of these datasets can be found in the given references and links.
@@
We follow the approach outlined in (Zhu et al., 2021). For the Avazu dataset, we transform the timestamp field it contains into three new feature fields: hour, weekday, and weekend. For the Criteo and KDD12 dataset, we discretize the numerical feature fields by rounding down each numeric value $x$ to $\lfloor\log^{2}(x)\rfloor$ if $x>2$, and $x=1$ otherwise. We set a threshold to replace infrequent categorical features with a default ”OOV” token. We set the threshold to 10 for Criteo, KKBox, and KDD12, 2 for Avazu and iPinYou, and 1 for the small dataset ML-1M. More specific data processing procedures and results can be found in our open-source run logs [^7] https://anonymous.4open.science/r/DCNv3-E352/checkpoints/ and configuration files, which we do not elaborate on here.
@@
We compared DCNv3 with some SOTA models (* denotes Integrating the original model with DNN networks): (1) Since ECN is a standalone network that performs explicit feature interactions, we compare it with several models that also perform explicit feature interactions on two large-scale datasets. For example, LR (2007) (Richardson et al., 2007) implements first-order feature interactions; FM and its derivative models FM (2010) (Rendle, 2010), FwFM (2018) (Pan et al., 2018), AFM (2017) (Xiao et al., 2017), FmFM (2021) (Sun et al., 2021) implement second-order feature interactions; and CrossNetv1 (2017) (Wang et al., 2017), CrossNetv2 (2021) (Wang et al., 2021b), CIN (2018) (Lian et al., 2018), AutoInt (2019) (Song et al., 2019), AFN (2020) (Cheng et al., 2020), FiGNN (2019) (Li et al., 2019) implement higher-order feature interactions. (2) To verify the superiority of DCNv3 over models that include implicit feature interactions, we further select several high-performance representative baselines, such as PNN (Qu et al., 2016) and Wide & Deep (Cheng et al., 2016) (2016); DeepFM (Guo et al., 2017) and DCNv1 (Wang et al., 2017) (2017); xDeepFM (2018) (Lian et al., 2018); AutoInt* (2019) (Song et al., 2019); AFN* (2020) (Cheng et al., 2020); DCNv2 (Wang et al., 2021b) and EDCN (Chen et al., 2021), MaskNet (Wang et al., 2021a) (2021); CL4CTR (Wang et al., 2023b), EulerNet (Tian et al., 2023), FinalMLP (Mao et al., 2023), and FINAL (Zhu et al., 2023b) (2023).
@@
We implement all models using PyTorch (Paszke et al., 2019) and refer to existing works (Zhu et al., 2021; Huawei, 2021). We employ the Adam optimizer (Kingma and Ba, 2014) to optimize all models, with a default learning rate set to 0.001. For the sake of fair comparison, we set the embedding dimension to 128 for KKBox and 16 for the other datasets (Zhu et al., 2021, 2022a). The batch size is set to 4,096 on the Criteo, ML-1M, and iPinYou datasets and 10,000 on the other datasets. To prevent overfitting, we employ early stopping with a patience value of 2. The hyperparameters of the baseline model are configured and fine-tuned based on the optimal values provided in (Huawei, 2021; Zhu et al., 2021) and their original paper. Further details on model hyperparameters and dataset configurations are available in our straightforward and accessible running logs [^8] https://anonymous.4open.science/r/DCNv3-E352/checkpoints/ and are not reiterated here.
@@ 4.2. Overall Performance (RQ1)
Higher-order CrossNetv1 (Wang et al., 2017) 0.3779 78.39 0.4456 80.59
@@
Table 4. Performance comparison of different deep CTR models. ”*”: Integrating the original model with DNN networks. Meanwhile, we conduct a two-tailed T-test to assess the statistical significance between our models and the best baseline ($\star$: $p$ 1e-3). Abs.Imp represents the absolute performance improvement of DCNv3 over the strongest baseline. Typically, CTR researchers consider an improvement of 0.001 (0.1%) in Logloss and AUC to be statistically significant (Wang et al., 2017; Chen et al., 2021; Wang et al., 2023b; Zhu et al., 2021).
  Avazu Criteo ML-1M KDD12 iPinYou KKBox
[…]
DNN (Covington et al., 2016) 0.3726 79.18 0.4380 81.40 0.3100 90.30 0.1502 80.52 0.005545 78.06 0.4811 85.01
PNN (Qu et al., 2016) 0.3719 79.32 0.4378 81.42 0.3070 90.42 0.1504 80.47 0.005544 78.13 0.4793 85.15
Wide & Deep (Cheng et al., 2016) 0.3725 79.20 0.4376 81.42 0.3056 90.45 0.1504 80.48 0.005542 78.09 0.4852 85.04
DeepFM (Guo et al., 2017) 0.3723 79.21 0.4375 81.43 0.3073 90.51 0.1501 80.60 0.005549 77.94 0.4785 85.31
DCNv1 (Wang et al., 2017) 0.3725 79.21 0.4376 81.44 0.3156 90.38 0.1501 80.59 0.005541 78.13 0.4766 85.31
xDeepFM (Lian et al., 2018) 0.3722 79.24 0.4376 81.43 0.3054 90.47 0.1501 80.62 0.005534 78.25 0.4772 85.35
AutoInt* (Song et al., 2019) 0.3722 79.24 0.4390 81.32 0.3112 90.45 0.1502 80.57 0.005544 78.16 0.4773 85.34
AFN* (Cheng et al., 2020) 0.3727 79.21 0.4384 81.38 0.3048 90.53 0.1499 80.70 0.005539 78.17 0.4842 84.89
DCNv2 (Wang et al., 2021b) 0.3724 79.22 0.4376 81.45 0.3098 90.56 0.1502 80.59 0.005539 78.26 0.4787 85.31
MaskNet (Wang et al., 2021a) 0.3716 79.36 0.4387 81.34 0.3080 90.34 0.1498 80.79 0.005556 77.85 0.5003 84.79
CL4CTR (Wang et al., 2023b) 0.3724 79.21 0.4383 81.35 0.3074 90.33 0.1502 80.56 0.005543 78.06 0.4972 83.78
FINAL (Zhu et al., 2023b) 0.3712 79.41 0.4371 81.49 0.3035 90.53 0.1498 80.74 0.005540 78.13 0.4795 85.08
DCNv3 0.3695 79.70 0.4358 81.62 0.3001 90.74 0.1494 80.98 0.005535 78.56 0.4765 85.57
@@
To further comprehensively investigate the performance superiority and generalization ability of DCNv3 on various CTR datasets (e.g., large-scale sparse datasets), we select 15 representative baseline models and 6 benchmark datasets. We highlight the performance of ECN and DCNv3 in bold and underline the best baseline performance. Table 4 presents the experimental results, from which we can make the following observations:
  • Overall, DCNv3 achieves the best performance across all six datasets, with an average AUC improvement of 0.21% over the strongest baseline model and an average Logloss decrease of 0.11%, both exceeding the statistically significant threshold of 0.1%. This demonstrates the effectiveness of DCNv3.
  • The FinalMLP model achieves good performance on the Avazu and Criteo datasets, surpassing most CTR models that combine explicit and implicit feature interactions. This demonstrates the effectiveness of implicit feature interactions. Consequently, most CTR models attempt to integrate DNN into explicit feature interaction models to enhance performance. However, DCNv3 achieves SOTA performance using only explicit feature interactions, indicating the effectiveness and potential of modeling with explicit feature interactions alone.
  • DCNv3 achieves performance improvements over ECN across all six datasets, demonstrating the effectiveness of LCN in capturing low-order feature interactions and the Tri-BCE loss. Notably, on the iPinYou dataset, we observe that all models have Logloss values around the 0.0055 level. This is due to the imbalance between positive and negative samples in the dataset (Lin et al., 2024), and other works have reported similar results (Qu et al., 2016, 2018).

### v7

@@ 3.3. Tri-BCE Loss
Tri-BCE is aimed at providing tailored supervision signals for different sub-networks. The calculation process and balancing method of the multi-loss are illustrated in Figure 4. We use the widely adopted binary cross-entropy loss (Wang et al., 2023a; Zhu et al., 2023b; Mao et al., 2023) (i.e., Logloss) as both the primary and auxiliary loss for the FCN:
@@
$\mathcal{L}_{\scriptscriptstyle\mathbb{D}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}\log\left(\hat{y}_{\scriptscriptstyle\mathbb{D},i}\right)+\left(1-y_{i}\right)\log\left(1-\hat{y}_{\scriptscriptstyle\mathbb{D},i}\right)\right),$
$\mathcal{L}_{\scriptscriptstyle\mathbb{S}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}\log\left(\hat{y}_{\scriptscriptstyle\mathbb{S},i}\right)+\left(1-y_{i}\right)\log\left(1-\hat{y}_{\scriptscriptstyle\mathbb{S},i}\right)\right),$
@@
where $y$ denotes the true labels, $N$ denotes the batch size, $\mathcal{L}_{\scriptscriptstyle\mathbb{D}}$ and $\mathcal{L}_{\scriptscriptstyle\mathbb{S}}$ represent the auxiliary losses for the prediction results of ECN and LCN, respectively, and $\mathcal{L}$ represents the primary loss. To provide each sub-network with suitable supervision signals, we assign them adaptive weights, $\mathbf{w}_{\scriptscriptstyle\mathbb{D}}=\texttt{max}(0,\mathcal{L}_{\scriptscriptstyle\mathbb{D}}-\mathcal{L})$ and $\mathbf{w}_{\scriptscriptstyle\mathbb{S}}=\texttt{max}(0,\mathcal{L}_{\scriptscriptstyle\mathbb{S}}-\mathcal{L})$, and jointly train them to achieve Tri-BCE loss:
(8) $\mathcal{L}_{\text{Tri}}$ $=\mathcal{L}+\mathbf{w}_{\scriptscriptstyle\mathbb{D}}\cdot\mathcal{L}_{\scriptscriptstyle\mathbb{D}}+\mathbf{w}_{\scriptscriptstyle\mathbb{S}}\cdot\mathcal{L}_{\scriptscriptstyle\mathbb{S}},$
@@
as demonstrated by (Li et al., 2024c), providing a single supervision signal to sub-networks is often suboptimal. Our proposed Tri-BCE loss helps sub-networks learn better parameters by providing adaptive weights that change throughout the learning process. Theoretically, we can derive the gradients obtained by $\hat{y}_{\scriptscriptstyle\mathbb{D}}$:
@@
(9) $\nabla_{(\hat{y}^{+}_{\scriptscriptstyle\mathbb{D}})}\mathcal{L}_{\text{Tri}}$ $=-\frac{1}{N}\cdot\frac{\partial\left(\log\hat{y}^{+}+\mathbf{w}_{\scriptscriptstyle\mathbb{D}}\log\hat{y}^{+}_{\scriptscriptstyle\mathbb{D}}\right)}{\partial\hat{y}^{+}_{\scriptscriptstyle\mathbb{D}}}=-\frac{1}{N}\left(\frac{1}{2\hat{y}^{+}}+\frac{\mathbf{w}_{\scriptscriptstyle\mathbb{D}}}{\hat{y}^{+}_{\scriptscriptstyle\mathbb{D}}}\right),$
@@
$\nabla_{(\hat{y}^{-}_{\scriptscriptstyle\mathbb{D}})}\mathcal{L}_{\text{Tri}}$ $=-\frac{1}{N}\cdot\frac{\partial\left(\log(1-\hat{y}^{-})+\mathbf{w}_{\scriptscriptstyle\mathbb{D}}\log(1-\hat{y}^{-}_{\scriptscriptstyle\mathbb{D}})\right)}{\partial\hat{y}^{-}_{\scriptscriptstyle\mathbb{D}}}$
$=\frac{1}{N}\left(\frac{1}{2(1-\hat{y}^{-})}+\frac{\mathbf{w}_{\scriptscriptstyle\mathbb{D}}}{1-\hat{y}^{-}_{\scriptscriptstyle\mathbb{D}}}\right),$
@@
where $\nabla_{(\hat{y}^{+}_{\scriptscriptstyle\mathbb{D}})}$ and $\nabla_{(\hat{y}^{-}_{\scriptscriptstyle\mathbb{D}})}$ represent the gradients received by $\hat{y}_{\scriptscriptstyle\mathbb{D}}$ for positive and negative samples, respectively. Similarly, the gradient signals received by $\hat{y}_{\scriptscriptstyle\mathbb{S}}$ are consistent with those of $\hat{y}_{\scriptscriptstyle\mathbb{D}}$, so we do not elaborate further. It can be observed that $\hat{y}_{\scriptscriptstyle\mathbb{D}}$ and $\hat{y}_{\scriptscriptstyle\mathbb{S}}$ both have the same gradient terms $\frac{1}{2\hat{y}^{+}}$ and $\frac{1}{2(1-\hat{y}^{-})}$, indicating that training both sub-networks with a single loss results in same supervision signals, which is detrimental to the model’s learning. However, our Tri-BCE loss additionally provides dynamically adjusted gradient terms based on $\mathbf{w}_{\scriptscriptstyle\mathbb{D}}$ and $\mathbf{w}_{\scriptscriptstyle\mathbb{S}}$, ensuring that the sub-networks are directly influenced by the true labels $y$ and adaptively adjust their weights according to the difference between the primary and auxiliary losses. Therefore, Tri-BCE loss provides the sub-networks with more suitable supervision signals.
@@ 3.4. Complexity Analysis
  Model Embedding Implicit interaction Explicit interaction
DCNv1 (Wang et al., 2017) O($dfs$) O($|W_{\Psi}|$) O($2DL$)
DCNv2 (Wang et al., 2021b) O($dfs$) O($|W_{\Psi}|$) O($D^{2}L$)
EDCN (Chen et al., 2021) O($dfs$) O($D^{2}L$) O($D^{2}L$)
GDCN (Wang et al., 2023a) O($dfs$) O($|W_{\Psi}|$) O($2D^{2}L$)
ECN O($dfs$) - O($\frac{D^{2}L}{2}$)
FCN O($dfs$) - O($D^{2}L$)
@@
  • Except for our proposed ECN and FCN, all other models include implicit interaction to enhance predictive performance, which incurs additional computational costs and reduces the overall interpretability of the model.
@@
  • Our FCN model uses the Tri-BCE loss function, which theoretically has a time complexity for loss computation three times higher than other models. However, in practical training, due to optimizations in parallel computation, its training cost is comparable to some models already deployed in production environments (e.g. FinalMLP(Mao et al., 2023), FINAL(Zhu et al., 2023b)) and does not reach the theoretical threefold increase. This is validated in Figure 5. Moreover, this design has no impact on the inference speed.
@@ 4. Experiments
In this section, we conduct comprehensive experiments on six CTR prediction datasets to validate the effectiveness, efficiency, and interpretability of FCN, and address the following research questions (RQs):
  • RQ1 Do FCN outperform other CTR models in terms of performance? Do they perform well on large-scale and highly sparse datasets?
  • RQ2 Are FCN more efficient compared to other CTR models?
  • RQ3 Do FCN possess interpretability and the ability to filter noise?
@@ 4.1. Experiment Setup
We evaluate FCN on six CTR prediction datasets: Avazu [^2] https://www.kaggle.com/c/avazu-ctr-prediction (Zhu et al., 2021), Criteo [^3] https://www.kaggle.com/c/criteo-display-ad-challenge (Zhu et al., 2021), ML-1M [^4] https://grouplens.org/datasets/movielens (Song et al., 2019), KDD12555https://www.kaggle.com/c/kddcup2012-track2 (Song et al., 2019), iPinYou [^6] https://contest.ipinyou.com/ (Qu et al., 2018), and KKBox [^7] https://www.kkbox.com/intl (Zhu et al., 2022a). Table 2 provides detailed information about these datasets. A more detailed description of these datasets can be found in the given references and links.
@@
We follow the approach outlined in (Zhu et al., 2021). For the Avazu dataset, we transform the timestamp field it contains into three new feature fields: hour, weekday, and weekend. For the Criteo and KDD12 dataset, we discretize the numerical feature fields by rounding down each numeric value $x$ to $\lfloor\log^{2}(x)\rfloor$ if $x>2$, and $x=1$ otherwise. We set a threshold to replace infrequent categorical features with a default ”OOV” token. We set the threshold to 10 for Criteo, KKBox, and KDD12, 2 for Avazu and iPinYou, and 1 for the small dataset ML-1M. More specific data processing procedures and results can be found in our open-source run logs [^8] https://github.com/salmon1802/FCN/tree/master/checkpoints/ and configuration files, which we do not elaborate on here.
@@
We compared FCN with some SOTA models (* denotes Integrating the original model with DNN networks): (1) Since ECN is a standalone network that performs explicit feature interactions, we compare it with several models that also perform explicit feature interactions on two large-scale datasets. For example, LR (2007) (Richardson et al., 2007) implements first-order feature interactions; FM and its derivative models FM (2010) (Rendle, 2010), FwFM (2018) (Pan et al., 2018), AFM (2017) (Xiao et al., 2017), FmFM (2021) (Sun et al., 2021) implement second-order feature interactions; and HOFM (2016) (Blondel et al., 2016), CrossNetv1 (2017) (Wang et al., 2017), CrossNetv2 (2021) (Wang et al., 2021b), CIN (2018) (Lian et al., 2018), AutoInt (2019) (Song et al., 2019), AFN (2020) (Cheng et al., 2020), FiGNN (2019) (Li et al., 2019) implement higher-order feature interactions. (2) To verify the superiority of FCN over models that include implicit feature interactions, we further select several high-performance representative baselines, such as PNN (Qu et al., 2016) and Wide & Deep (Cheng et al., 2016) (2016); DeepFM (Guo et al., 2017) and DCNv1 (Wang et al., 2017) (2017); xDeepFM (2018) (Lian et al., 2018); AutoInt* (2019) (Song et al., 2019); AFN* (2020) (Cheng et al., 2020); DCNv2 (Wang et al., 2021b) and EDCN (Chen et al., 2021), MaskNet (Wang et al., 2021a) (2021); CL4CTR (Wang et al., 2023b), EulerNet (Tian et al., 2023), FinalMLP (Mao et al., 2023), FINAL (Zhu et al., 2023b) (2023), and RFM (Tian et al., 2024) (2024).
@@
We implement all models using PyTorch (Paszke et al., 2019) and refer to existing works (Zhu et al., 2021; Huawei, 2021). We employ the Adam optimizer (Kingma and Ba, 2014) to optimize all models, with a default learning rate set to 0.001. For the sake of fair comparison, we set the embedding dimension to 128 for KKBox and 16 for the other datasets (Zhu et al., 2021, 2022a). The batch size is set to 4,096 on the Criteo, ML-1M, and iPinYou datasets and 10,000 on the other datasets. To prevent overfitting, we employ early stopping with a patience value of 2. The hyperparameters of the baseline model are configured and fine-tuned based on the optimal values provided in (Huawei, 2021; Zhu et al., 2021) and their original paper. Further details on model hyperparameters and dataset configurations are available in our straightforward and accessible running logs8 and are not reiterated here.
@@ 4.2. Overall Performance (RQ1)
Higher-order HOFM (Blondel et al., 2016) 0.3754 78.91 0.4411 81.07
CrossNetv1 (Wang et al., 2017) 0.3779 78.39 0.4456 80.59
@@
Table 4. Performance comparison of different deep CTR models. ”*”: Integrating the original model with DNN networks. Meanwhile, we conduct a two-tailed T-test to assess the statistical significance between our models and the best baseline ($\star$: $p$ 1e-3). Abs.Imp represents the absolute performance improvement of FCN over the strongest baseline. Typically, CTR researchers consider an improvement of 0.001 (0.1%) in Logloss and AUC to be statistically significant (Wang et al., 2017; Chen et al., 2021; Wang et al., 2023b; Zhu et al., 2021).
  Avazu Criteo ML-1M KDD12 iPinYou KKBox
[…]
DNN (Covington et al., 2016) 0.3721 79.27 0.4380 81.40 0.3100 90.30 0.1502 80.52 0.005545 78.06 0.4811 85.01
PNN (Qu et al., 2016) 0.3712 79.44 0.4378 81.42 0.3070 90.42 0.1504 80.47 0.005544 78.13 0.4793 85.15
Wide & Deep (Cheng et al., 2016) 0.3720 79.29 0.4376 81.42 0.3056 90.45 0.1504 80.48 0.005542 78.09 0.4852 85.04
DeepFM (Guo et al., 2017) 0.3719 79.30 0.4375 81.43 0.3073 90.51 0.1501 80.60 0.005549 77.94 0.4785 85.31
DCNv1 (Wang et al., 2017) 0.3719 79.31 0.4376 81.44 0.3156 90.38 0.1501 80.59 0.005541 78.13 0.4766 85.31
xDeepFM (Lian et al., 2018) 0.3718 79.33 0.4376 81.43 0.3054 90.47 0.1501 80.62 0.005534 78.25 0.4772 85.35
AutoInt* (Song et al., 2019) 0.3746 79.02 0.4390 81.32 0.3112 90.45 0.1502 80.57 0.005544 78.16 0.4773 85.34
AFN* (Cheng et al., 2020) 0.3726 79.29 0.4384 81.38 0.3048 90.53 0.1499 80.70 0.005539 78.17 0.4842 84.89
DCNv2 (Wang et al., 2021b) 0.3718 79.31 0.4376 81.45 0.3098 90.56 0.1502 80.59 0.005539 78.26 0.4787 85.31
MaskNet (Wang et al., 2021a) 0.3711 79.43 0.4387 81.34 0.3080 90.34 0.1498 80.79 0.005556 77.85 0.5003 84.79
CL4CTR (Wang et al., 2023b) 0.3722 79.25 0.4383 81.35 0.3074 90.33 0.1502 80.56 0.005543 78.06 0.4972 83.78
FINAL (Zhu et al., 2023b) 0.3712 79.41 0.4371 81.49 0.3035 90.53 0.1498 80.74 0.005540 78.13 0.4800 85.14
RFM (Tian et al., 2024) 0.3723 79.24 0.4374 81.47 0.3048 90.51 0.1506 80.73 0.005540 78.25 0.4853 84.70
FCN 0.3695 79.70 0.4358 81.62 0.3001 90.74 0.1494 80.98 0.005535 78.56 0.4765 85.57
@@
To further comprehensively investigate the performance superiority and generalization ability of FCN on various CTR datasets (e.g., large-scale sparse datasets), we select 15 representative baseline models and 6 benchmark datasets. We highlight the performance of ECN and FCN in bold and underline the best baseline performance. Table 4 presents the experimental results, from which we can make the following observations:
  • Overall, FCN achieves the best performance across all six datasets, with an average AUC improvement of 0.21% over the strongest baseline model and an average Logloss decrease of 0.11%, both exceeding the statistically significant threshold of 0.1%. This demonstrates the effectiveness of FCN.
  • The FinalMLP model achieves good performance on the Avazu and Criteo datasets, surpassing most CTR models that combine explicit and implicit feature interactions. This demonstrates the effectiveness of implicit feature interactions. Consequently, most CTR models attempt to integrate DNN into explicit feature interaction models to enhance performance. However, FCN achieves SOTA performance using only explicit feature interactions, indicating the effectiveness and potential of modeling with explicit feature interactions alone.
  • FCN achieves performance improvements over ECN across all six datasets, demonstrating the effectiveness of LCN in capturing low-order feature interactions and the Tri-BCE loss. Notably, on the iPinYou dataset, we observe that all models have Logloss values around the 0.0055 level. This is due to the imbalance between positive and negative samples in the dataset (Lin et al., 2024), and other works have reported similar results (Qu et al., 2016, 2018).